OpenStudy (nobody):

What is the value of y in the equation 4(5y − 8 − 2) = 185 − 15? 8.5 9.7 10.5 42.5

OpenStudy (will.h):

4(5y − 8 − 2) = 185 − 15 distribute 4 to the terms insides the parentheses 20y - 32 - 8 = 185 - 15 Simplify 20y -40 = 170 20y = 210 y = 10.5

OpenStudy (tarnett1):

20y-32-8+=170 20y-24=170 then you subtract add 24 to each side the answer is 9.7

OpenStudy (nobody):

wow you are amazing! Thank you so much! @Will.H

OpenStudy (will.h):

my pleasure :)

OpenStudy (nobody):

I got a 100%!! @Will.H

OpenStudy (will.h):

Congrats mate :)